OpenStudy (anonymous):

if y=sin^-1(5x), then dy/dx=

OpenStudy (anonymous):

chain rule for this one. but first you need to know the derivative of \[\sin^{-1}(x)\] do you know it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is \[\frac{d}{dx}\sin^{-1}(x)=\frac{1}{\sqrt{1-x^2}}\] then use the chain rule

OpenStudy (anonymous):

so \[\frac{d}{dx}\sin^{-1}(5x)=\frac{5}{\sqrt{1-(5x)^2}}\]

OpenStudy (anonymous):

but the answer is 5\square root(1-25x^2) but idk how thats the answer

OpenStudy (anonymous):